# Geological Everything Model 3D: A Promptable Foundation Model for Unified and Zero-Shot Subsurface Understanding

### v2

@@ 2.5.2 Acoustic impedance modeling on the Teapot Dome
In Figure 5 (c), GEM demonstrates superior performance in capturing thinlayer structures with greater clarity and continuity. It effectively suppresses the vertical ringing and lateral instability commonly observed in traditional methods (Figure 11). While maintaining consistency with the seismic structural framework, the impedance volumes generated by GEM exhibit higher spatial resolution and stronger stratigraphic continuity, particularly in zones with thin interbedded sandstone and mudstone. Figure 5 (c-3) presents the blind-well test results, demonstrating a high degree of alignment between GEM’s predictions and the impedance trends observed in well logs. The corresponding MAE values further confirm the model’s accuracy. In addition, Appendices Figure 11 (a) provides a qualitative comparison of property modeling results between GEM and several baseline methods. GEM consistently outperforms the others, particularly in terms of vertical resolution and the preservation of geological structures.
2.5.3 Multi-property modeling on the Netherlands F3
The Netherlands F3 dataset is a widely used benchmark for 3D seismic interpretation and property modeling. Geologically, it features multi-phase fault systems, progradational slope deposits, and a regionally extensive angular unconformity that separates genetically distinct stratigraphic packages. These structures present strong phase shifts, impedance discontinuities, and highamplitude reflectors—making property modeling in this dataset particularly challenging. Such complex features demand models capable of capturing nonlinear responses and maintaining stratigraphic coherence across discontinuities.
@@ 2.5.3 Multi-property modeling on the Netherlands F3
2.5.4 Multi-property modeling on the Delft
The Delft dataset presents a particularly challenging scenario for property modeling due to its complex depositional and tectonic history. The region is dominated by fluvial facies deposits with strong lateral heterogeneity, and is further complicated by igneous intrusion from depth, which have deformed overlying strata through structural folding and reverse faulting. These features disrupt stratigraphic continuity and create irregular distributions of physical properties, making it difficult for conventional methods to generate models that are both geologically plausible and spatially coherent.
@@ 3.1 Unified Perspective and Overall Architecture
We propose a unified view in which subsurface structural and stratigraphic interpretation, geobody segmentation, and property modeling can treated as tasks operating at different levels of abstraction of a common geologic framework. Based on this insight, we designed GEM to unify these diverse tasks through a prompt-driven generative process that propagates interactive sparse prompts along structural features inferred from subsurface imaging data. By leveraging shared structural constraints—including stratigraphic continuity, fault systems, regional geobody characteristics, stratigraphic hierarchy and spatial context relationships—the model enables coherent task integration and collaborative optimization within a single framework.
@@
To realize this unified modeling strategy, GEM employs a two-stage training framework consisting of self-supervised pretraining followed by supervised fine-tuning with mixed prompts and labels drawn from diverse tasks (Figure 1 (b, c)). The pretraining stage employs a masked image modeling strategy to learn generalizable structural representations from large-scale real and synthetic seismic data This stage establishes a foundational feature representation for downstream tasks by learning the broad geological characteristics present in real data, thereby accelerating the convergence of subsequent task-specific training. The fine-tuning stage supervises conditional generation with a mixture of task-specific prompts and labels from diverse interpretation tasks, implemented under a Generative Adversarial Network (GAN) framework [61].
@@ 3.2 Self Supervised Pre-training
The training dataset comprises a large collection of both field-acquired and synthetic seismic volumes. These volumes are randomly cropped into cubes of size $160\times 160\times 160$. Pretraining is conducted on 8×8$\times$8 × NVIDIA H20 GPUs with a batch size of 64, using the AdamW optimizer. The learning rate is set to $0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and the training proceeds for 500,000 iterations.
@@ 3.3 Structure-aware Perceptual Network
Training of the perceptual network is conducted on 8×8$\times$8 × NVIDIA H20 GPUs with a batch size of $96$, using the AdamW optimizer with a learning rate of $0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a total of 300,000 iterations.
@@ 3.4 Supervised Fine-tuning with Mixed Prompting
Due to the ambiguity and anisotropy inherent in subsurface imaging data, which often result in non-uniqueness, practical applications tend to prioritize the reliability of a solution among all possible outcomes rather than its absolute correctness. This consideration is particularly critical in tasks such as structural modeling, RGT estimation, and property modeling, where the objective is to recover geologically plausible structures from incomplete information. Under such circumstances, only relative correctness can be meaningfully defined. To address this challenge, we adopt a relativistic GAN strategy [64]. This approach shifts the modeling objective away from generating a single correct solution and instead encourages the model to learn how to select relatively better and geologically more credible structures from among multiple plausible solutions. This modeling paradigm aligns more naturally with the inherent uncertainty and non-uniqueness of seismic data. The corresponding formulation is given as follows,
@@
Training was performed on 8×8$\times$8 × Nvidia H20 GPUs with a batch size of 8 for a total of 500,000 iterations. The first 300,000 iterations used an input size of $320\times 400\times 128$ (timeline, inline, crossline), followed by 200,000 iterations with an increased size of $512\times 512\times 128$ under PyTorch’s checkpoint mode. During training, the inline and crossline dimensions were randomly transposed to improve spatial generalization. The Adam optimizer is employed with the Two Time-Scale Update Rule (TTUR) [65]. The learning rate is set to 0.001 for the generator and 0.004 for the discriminator, with $\beta_{1}=0$ and $\beta_{2}=0.9$.
@@ 4 Conclusion
We introduce GEM, a prompt-based foundation model for subsurface imaging that unifies structural interpretation, stratigraphic analysis, geobody segmentation, and property modeling within a single generative framework. By reformulating all these tasks as prompt-guided conditional generation over structurally inferred latent spaces, GEM supports flexible human interaction and generalizes across surveys, data types, and geological settings without task-specific retraining.
Unlike conventional models constrained to specific tasks and surveys, GEM leverages large-scale self-supervised pretraining and adversarial fine-tuning with heterogeneous prompts to achieve zero-shot generalization across a broad range of subsurface interpretation and modeling tasks. It demonstrates strong performance on real-world seismic and radar datasets, including the delineation of complex fault systems, Martian polar stratigraphy, full-volume seismic stratigraphy, diverse geobody segmentation, and multi-property modeling guided by sparse well-log inputs. By integrating multimodal expert cues with structural priors, GEM serves as a new interface between geoscientific expertise and generative, prompt-driven reasoning in Earth science.
To our knowledge, GEM is the first generative foundation model in subsurface understanding that supports expert-in-the-loop interaction, prompt-guided reasoning, and zero-shot generalization across tasks, surveys, and data types. More broadly, this work establishes a new modeling paradigm for scientific AI in domains where data is structurally rich but semantically sparse, and where insight emerges from contextual, expert-informed interpretation. GEM exemplifies how generative models can be steered by structural priors, enabling interpretable, scalable, and collaborative modeling of the Earth’s interior. We envision this structure-aware, promptable approach extending to broader domains in scientific machine learning, including physical simulation, inverse problems, and multi-modal Earth system modeling.
[…]
@@ D.3 Property Modeling
Figure 11: Qualitative comparison of property modeling across diverse geological settings. (a, b) Acoustic impedance modeling on the Teapot Dome and F3 datasets using sparse well logs. Conventional AI models (UNet, DeepLabV3, HRNet) trained with task-specific supervision exhibit discontinuities and lateral instability, particularly on the F3 dataset with limited well coverage. In contrast, GEM produces laterally consistent impedance volumes using well logs as prompts, without task-specific retraining. (c) DTCO modeling in the structurally complex Poseidon dataset. Despite strong lateral variation and sparse well control, GEM captures fault displacements and unconformity terminations with high geometric fidelity. (d) Gamma ray modeling in the same region. These results highlight GEM’s robustness in property modeling, demonstrating strong structural awareness, lateral continuity, and adaptability across structurally complex surveys with limited well-log prompts.
Figure 11 presents the impedance modeling results of conventional methods on the Teapot and F3 datasets. These traditional approaches uniformly adopt a multidimensional modeling strategy, wherein sparse 1D well-log labels are used to supervise 3D seismic volumes. Compared to the property modeling outcomes generated by GEM, the results from conventional impedance modeling exhibit insufficient lateral resolution and pronounced discontinuities along the reflectors. This issue is especially evident in the F3 dataset, where numerous anomalous value regions appear. The root cause lies in the fact that only four wells are available in the F3 block, with just three used for training, leading to severe overfitting and a lack of reliability in the network’s predictions.
@@
As shown in Figure 11 (c-1), the DTCO well log exhibits strong lateral variations in its upper section. Combined with the complex lower fault system and sparse well coverage, these factors make modeling particularly difficult. Nevertheless, GEM is still able to produce reliable property modeling results. In (c-2), the predicted property volume reveals a fault system consistent with seismic imaging, clearly delineating the vertical termination of unconformities and their superposition on fault structures, thereby revealing the deformation history of stratums under tectonic processes. The predicted property responses exhibit strong geometric continuity along fault trends and accurately capture intense heterogeneity in fault intersection zones.
@@
This example demonstrates that GEM maintains strong structural awareness and expressive capacity in property modeling even under sparse well control. The results further validate the model’s robustness in parsing structural frameworks under prompt guidance, highlighting its generalization ability and practical potential in geologically complex and data-limited scenarios.
@@ Appendix F Training and Inference Cost
For interactive inference tasks such as structural interpretation and geobody segmentation, we recommend resizing the data to lower resolutions or smaller volumes on the GPU to improve interactivity. In contrast, tasks like property modeling and RGT estimation benefit from inference at the original resolution or volume size, which can be efficiently handled on CPU memory (CPU memory is more accessible and easier to scale, allowing for inference on larger volumes). This sacrifices inference speed but yields higher voxel-level accuracy and resolution.

### v3

@@ 2.5.2 Acoustic impedance modeling on the Teapot Dome
In Figure 5 (c), GEM demonstrates superior performance in capturing thinlayer structures with greater clarity and continuity. It effectively suppresses the vertical ringing and lateral instability commonly observed in traditional methods (Figure 11). While maintaining consistency with the seismic structural framework, the impedance volumes generated by GEM exhibit higher spatial resolution and stronger stratigraphic continuity, particularly in zones with thin interbedded sandstone and mudstone. Figure 5 (c-3) presents the blind-well test results, demonstrating a high degree of alignment between GEM’s predictions and the impedance trends observed in well logs. The corresponding MAE values further confirm the model’s accuracy. In addition, Appendices Figure 11 (a) provides a qualitative comparison of physical property modeling results between GEM and several baseline methods. GEM consistently outperforms the others, particularly in terms of vertical resolution and the preservation of geological structures.
2.5.3 Multi-physical-property modeling on the Netherlands F3
The Netherlands F3 dataset is a widely used benchmark for 3D seismic interpretation and physical property modeling. Geologically, it features multi-phase fault systems, progradational slope deposits, and a regionally extensive angular unconformity that separates genetically distinct stratigraphic packages. These structures present strong phase shifts, impedance discontinuities, and highamplitude reflectors—making physical property modeling in this dataset particularly challenging. Such complex features demand models capable of capturing nonlinear responses and maintaining stratigraphic coherence across discontinuities.
@@ 2.5.3 Multi-physical-property modeling on the Netherlands F3
2.5.4 Multi-physical-property modeling on the Delft
The Delft dataset presents a particularly challenging scenario for physical property modeling due to its complex depositional and tectonic history. The region is dominated by fluvial facies deposits with strong lateral heterogeneity, and is further complicated by igneous intrusion from depth, which have deformed overlying strata through structural folding and reverse faulting. These features disrupt stratigraphic continuity and create irregular distributions of physical properties, making it difficult for conventional methods to generate models that are both geologically plausible and spatially coherent.
@@ 3.1 Unified Perspective and Overall Architecture
We propose a unified view in which subsurface structural and stratigraphic interpretation, geobody segmentation, and physical property modeling can treated as tasks operating at different levels of abstraction of a common geologic framework. Based on this insight, we designed GEM to unify these diverse tasks through a prompt-driven generative process that propagates interactive sparse prompts along structural features inferred from subsurface imaging data. By leveraging shared structural constraints—including stratigraphic continuity, fault systems, regional geobody characteristics, stratigraphic hierarchy and spatial context relationships—the model enables coherent task integration and collaborative optimization within a single framework.
@@
To realize this unified modeling strategy, GEM employs a two-stage training framework consisting of self-supervised pretraining followed by supervised fine-tuning with mixed prompts and labels drawn from diverse tasks (Figure 1 (b, c)). The pretraining stage employs a masked image modeling strategy to learn generalizable structural representations from large-scale real and synthetic seismic data This stage establishes a foundational feature representation for downstream tasks by learning the broad geological characteristics present in real data, thereby accelerating the convergence of subsequent task-specific training. The fine-tuning stage supervises conditional generation with a mixture of task-specific prompts and labels from diverse interpretation tasks, implemented under a Generative Adversarial Network (GAN) framework [61]. Together these design choices instantiate a physics informed deep learning paradigm in which geological constraints are injected as differentiable, structure aware guidance.
@@ 3.2 Self Supervised Pre-training
The training dataset comprises a large collection of both field-acquired and synthetic seismic volumes. These volumes are randomly cropped into cubes of size $160\times 160\times 160$. Pretraining is conducted on $8\times$ NVIDIA H20 GPUs with a batch size of 64, using the AdamW optimizer. The learning rate is set to $0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and the training proceeds for 500,000 iterations.
@@ 3.3 Structure-aware Perceptual Network
Training of the perceptual network is conducted on $8\times$ NVIDIA H20 GPUs with a batch size of $96$, using the AdamW optimizer with a learning rate of $0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a total of 300,000 iterations.
@@ 3.4 Supervised Fine-tuning with Mixed Prompting
Due to the ambiguity and anisotropy inherent in subsurface imaging data, which often result in non-uniqueness, practical applications tend to prioritize the reliability of a solution among all possible outcomes rather than its absolute correctness. This consideration is particularly critical in tasks such as structural modeling, RGT estimation, and physical property modeling, where the objective is to recover geologically plausible structures from incomplete information. Under such circumstances, only relative correctness can be meaningfully defined. To address this challenge, we adopt a relativistic GAN strategy [64]. This approach shifts the modeling objective away from generating a single correct solution and instead encourages the model to learn how to select relatively better and geologically more credible structures from among multiple plausible solutions. This modeling paradigm aligns more naturally with the inherent uncertainty and non-uniqueness of seismic data. The corresponding formulation is given as follows,
@@
The combination of the structure aware perceptual loss and relativistic adversarial training operationalizes physics informed objectives that promote geologically consistent and reliable predictions under limited supervision.
Training was performed on $8\times$ Nvidia H20 GPUs with a batch size of 8 for a total of 500,000 iterations. The first 300,000 iterations used an input size of $320\times 400\times 128$ (timeline, inline, crossline), followed by 200,000 iterations with an increased size of $512\times 512\times 128$ under PyTorch’s checkpoint mode. During training, the inline and crossline dimensions were randomly transposed to improve spatial generalization. The Adam optimizer is employed with the Two Time-Scale Update Rule (TTUR) [65]. The learning rate is set to 0.001 for the generator and 0.004 for the discriminator, with $\beta_{1}=0$ and $\beta_{2}=0.9$.
@@ 4 Conclusion
We introduce GEM, a prompt-based foundation model for subsurface imaging that unifies structural interpretation, stratigraphic analysis, geobody segmentation, and physical property modeling within a single generative framework. By reformulating all these tasks as prompt-guided conditional generation over structurally inferred latent spaces, GEM supports flexible human interaction and generalizes across surveys, data types, and geological settings without task-specific retraining.
Unlike conventional models constrained to specific tasks and surveys, GEM leverages large-scale self-supervised pretraining and adversarial fine-tuning with heterogeneous prompts to achieve zero-shot generalization across a broad range of subsurface interpretation and modeling tasks. It demonstrates strong performance on real-world seismic and radar datasets, including the delineation of complex fault systems, Martian polar stratigraphy, full-volume seismic stratigraphy, diverse geobody segmentation, and multi-physical-property modeling guided by sparse well-log inputs. By integrating multimodal expert cues with structural priors, GEM serves as a new interface between geoscientific expertise and generative, prompt-driven reasoning in Earth science.
To our knowledge, GEM is the first generative foundation model in subsurface understanding that supports expert-in-the-loop interaction, prompt-guided reasoning, and zero-shot generalization across tasks, surveys, and data types. More broadly, this work establishes a new modeling paradigm for scientific AI in domains where data is structurally rich but semantically sparse, and where insight emerges from contextual, expert-informed interpretation. GEM exemplifies how generative models can be steered by structural priors, enabling interpretable, scalable, and collaborative modeling of the Earth’s interior. We envision this structure-aware, promptable approach extending to broader domains in scientific machine learning, including physical simulation, inverse problems, and multi-modal Earth system modeling. In this sense, GEM advances the aims of physics informed machine learning by enabling data efficient and trustworthy predictive modeling and by supporting data driven scientific discovery in geophysics.
@@ D.3 Property Modeling
Figure 11: Qualitative comparison of physical property modeling across diverse geological settings. (a, b) Acoustic impedance modeling on the Teapot Dome and F3 datasets using sparse well logs. Conventional AI models (UNet, DeepLabV3, HRNet) trained with task-specific supervision exhibit discontinuities and lateral instability, particularly on the F3 dataset with limited well coverage. In contrast, GEM produces laterally consistent impedance volumes using well logs as prompts, without task-specific retraining. (c) DTCO modeling in the structurally complex Poseidon dataset. Despite strong lateral variation and sparse well control, GEM captures fault displacements and unconformity terminations with high geometric fidelity. (d) Gamma ray modeling in the same region. These results highlight GEM’s robustness in physical property modeling, demonstrating strong structural awareness, lateral continuity, and adaptability across structurally complex surveys with limited well-log prompts.
Figure 11 presents the impedance modeling results of conventional methods on the Teapot and F3 datasets. These traditional approaches uniformly adopt a multidimensional modeling strategy, wherein sparse 1D well-log labels are used to supervise 3D seismic volumes. Compared to the physical property modeling outcomes generated by GEM, the results from conventional impedance modeling exhibit insufficient lateral resolution and pronounced discontinuities along the reflectors. This issue is especially evident in the F3 dataset, where numerous anomalous value regions appear. The root cause lies in the fact that only four wells are available in the F3 block, with just three used for training, leading to severe overfitting and a lack of reliability in the network’s predictions.
@@
As shown in Figure 11 (c-1), the DTCO well log exhibits strong lateral variations in its upper section. Combined with the complex lower fault system and sparse well coverage, these factors make modeling particularly difficult. Nevertheless, GEM is still able to produce reliable physical property modeling results. In (c-2), the predicted property volume reveals a fault system consistent with seismic imaging, clearly delineating the vertical termination of unconformities and their superposition on fault structures, thereby revealing the deformation history of stratums under tectonic processes. The predicted property responses exhibit strong geometric continuity along fault trends and accurately capture intense heterogeneity in fault intersection zones.
@@
This example demonstrates that GEM maintains strong structural awareness and expressive capacity in physical property modeling even under sparse well control. The results further validate the model’s robustness in parsing structural frameworks under prompt guidance, highlighting its generalization ability and practical potential in geologically complex and data-limited scenarios.
@@ Appendix F Training and Inference Cost
For interactive inference tasks such as structural interpretation and geobody segmentation, we recommend resizing the data to lower resolutions or smaller volumes on the GPU to improve interactivity. In contrast, tasks like physical property modeling and RGT estimation benefit from inference at the original resolution or volume size, which can be efficiently handled on CPU memory (CPU memory is more accessible and easier to scale, allowing for inference on larger volumes). This sacrifices inference speed but yields higher voxel-level accuracy and resolution.